OpenStudy (anonymous):

please help me with this problem 29.78 the (.) showd to be a times that should be disrtibute

OpenStudy (anonymous):

\[29\times 78\] ?

OpenStudy (anonymous):

yeah just like that

OpenStudy (anonymous):

no calculator?

OpenStudy (anonymous):

it dont have to use calculatr it has to be distribute

OpenStudy (anonymous):

multiply using distibutive property

OpenStudy (anonymous):

\(9\times 8=72\) \(70\times 9=630\) \(20\times 8=160\) \(20\times 70=1400\) and now all that is left is to add

OpenStudy (anonymous):

ow okie thank u

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

other one XDD 58x31 it has to be distrbute too :">